# Relevance for Human Robot Collaboration

### v1

@@ Abstract
Effective human-robot collaboration (HRC) requires the robots to possess human-like intelligence. Inspired by the human’s cognitive ability to selectively process and filter elements in complex environments, this paper introduces a novel concept and scene-understanding approach termed ‘relevance.’ It identifies relevant components in a scene. To accurately and efficiently quantify relevance, we developed an event-based framework that selectively triggers relevance determination, along with a probabilistic methodology built on a structured scene representation. Simulation results demonstrate that the relevance framework and methodology accurately predict the relevance of a general HRC setup, achieving a precision of 0.99 and a recall of 0.94. Relevance can be broadly applied to several areas in HRC to improve task planning time by 79.56% compared with pure planning for a cereal task, reduce perception latency by up to 26.53% for an object detector, improve HRC safety by up to 13.50% and reduce the number of inquiries for HRC by 75.36%. A real-world demonstration showcases the relevance framework’s ability to intelligently assist humans in everyday tasks.
@@ V-C Relevance in HRC
One benefit of relevance in HRC is that relevance can help the robot better reason about the scene and the human’s requirements in a well-structured manner, which resembles how humans interact with other humans. This benefit can be demonstrated with a simple robot inquiry example in a coffee serving setup. The human-like robot will first detect the relevant classes and types of elements. If the relevance of any classes or types of elements is high, those classes or types are deemed necessary (coffee in this domain). Other relevant classes or types of elements are optional, and the robot asks the human if they need those relevant elements. The number of inquiries required using different methods is shown in Table I. It is shown that with precise and proper reasoning about the relevance of an element to a task or objective, the number of inquiries required for appropriate assistance is reduced by 75.36% compared with the number without relevance. Thus, human-robot interactions can be much more natural and fluent.
@@ VII Conclusion
This paper presents a novel concept and scene-understanding approach, termed ’relevance,’ designed to identify pertinent elements within a scene to better support human objectives. To quantify relevance, we developed an innovative framework that continuously monitors the scene and selectively activates relevance assessment. We developed a probabilistic methodology with a novel scene representation to efficiently and accurately quantify relevance. Simulations demonstrate the framework’s effectiveness, achieving a precision of 0.99 and a recall of 0.94. to efficiently and accurately determine relevance. Simulation results demonstrate the effectiveness of this framework and methodology, achieving a precision of 0.99 and a recall of 0.94. Relevance can significantly enhance HRC by improving task planning time by 79.56% compared to pure planning for a cereal task, reducing perception latency by up to 26.53% for object detection, enhancing safety by up to 13.50%, and decreasing the number of inquiries by 75.36%. The experimental results show a significant advancement toward creating a versatile robotic assistant capable of providing intelligent support to humans.

### v2

@@ Abstract
Effective human-robot collaboration (HRC) requires the robots to possess human-like intelligence. Inspired by the human’s cognitive ability to selectively process and filter elements in complex environments, this paper introduces a novel concept and scene-understanding approach termed ‘relevance.’ It identifies relevant components in a scene. To accurately and efficiently quantify relevance, we developed an event-based framework that selectively triggers relevance determination, along with a probabilistic methodology built on a structured scene representation. Simulation results demonstrate that the relevance framework and methodology accurately predict the relevance of a general HRC setup, achieving a precision of 0.99 and a recall of 0.94. Relevance can be broadly applied to several areas in HRC to improve task planning time by 79.56% compared with pure planning for a cereal task, reduce perception latency by up to 26.53% for an object detector, improve HRC safety by up to 13.50% and reduce the number of inquiries for HRC by 80.84%. A real-world demonstration showcases the relevance framework’s ability to intelligently assist humans in everyday tasks.
@@ V-C Relevance in HRC
One benefit of relevance in HRC is that relevance can help the robot better reason about the scene and the human’s requirements in a well-structured manner, which resembles how humans interact with other humans. This benefit can be demonstrated with a simple robot inquiry example in a coffee serving setup. The human-like robot will first detect the relevant classes and types of elements. If the relevance of any classes or types of elements is high, those classes or types are deemed necessary (coffee in this domain). Other relevant classes or types of elements are optional, and the robot asks the human if they need those relevant elements. The number of inquiries required using different methods is shown in Table I. It is shown that with precise and proper reasoning about the relevance of an element to a task or objective, the number of inquiries required for appropriate assistance is reduced by 80.84% compared with the number without relevance. Thus, human-robot interactions can be much more natural and fluent.
@@ VII Conclusion
This paper presents a novel concept and scene-understanding approach, termed ‘relevance,’ designed to identify pertinent elements within a scene to better support human objectives. To quantify relevance, we developed an innovative framework that continuously monitors the scene and selectively activates relevance assessment. We developed a probabilistic methodology with a novel scene representation to efficiently and accurately quantify relevance. Simulations demonstrate the framework’s effectiveness, achieving a precision of 0.99 and a recall of 0.94. to efficiently and accurately determine relevance. Relevance can significantly enhance HRC by improving task planning time by 79.56% compared to pure planning for a cereal task, reducing perception latency by up to 26.53% for object detection, enhancing safety by up to 13.50%, and decreasing the number of inquiries by 80.84%. The experimental results show a significant advancement toward creating a versatile robotic assistant capable of providing intelligent support to humans.